OpenStudy (hba):

A body traveling at constant speed of 2 km/h moves along a circular curve of radius 1000 m. Its acceleration will be A) 2 km/h² B) 3 km/h² C) 3.5 km/h² D) 4 km/h²

OpenStudy (anonymous):

D

OpenStudy (hba):

Can i know how did you do it ?

OpenStudy (anonymous):

the centripetal force acting on a body moving in a circle is mv^2/r ...equating it to the general force F=ma, you get a=v^2/r.....So, a= (4000m/s)^2 / 1000m = 4m/s^2

OpenStudy (anonymous):

iam sorry for wrong simplification...a= v^2/r = 2^2/1 = 4

OpenStudy (anonymous):

1000m = 1km

OpenStudy (eyust707):

I like to derive uniform circular motion from a circle.

OpenStudy (anonymous):

Did you get it?